Suppose that there are two very large reservoirs of water, one at a temperature of 89.0 °C and one at a temperature of 13.0 °C. These reservoirs are brought into thermal contact long enough for 42530 J of heat to flow from the hot water to the cold water. Assume that the reservoirs are large enough so that the temperatures do not change significantly. What is the total change in entropy resulting from this heat exchange between the hot water and the cold water? total change in entropy: 32 Calculate the amount of energy made unavailable for work by this increase in entropy. unt of energy unavailable for work: How much work could a Carnot engine do if it took in the given amount of heat (42530 J) from the hot water reservoir and exhausted heat to the cold water reservoir? work done by a Carnot engine:
Answer #1

A)

Entropy change, dS = (Q/Tl) + (-Q/Th)

dS = (42530/286) - (42530/362) = 31.22

B)

Unavailable work, E = 286 x 31.22 = 8928.95 J

C)

Work, W = 42530 (1 - (286/362)) = 8928.95 J

Comment in case any doubt please rate my answer....